OpenStudy (anonymous):

A 3-ft.-wide sidewalk surrounds a rectangular plot that measures 20 ft. by 30 ft. Find the area of the sidewalk.

OpenStudy (anonymous):

I have only these three choices for answer 936 ft.² 600 ft.² 336 ft.²

jimthompson5910 (jim_thompson5910):

Area of Plot+Sidewalk: A = LW A = (30+2*3)*(20+2*3) A = 936 ------------------------------------ Area of Plot only A = LW A = 30*20 A = 600 ------------------------------------ Area of Sidewalk only = (Area of Plot+Sidewalk) - (Area of Plot only) Area of Sidewalk only = 936-600 Area of Sidewalk only = 336 So the area of the sidewalk is 336 square feet

OpenStudy (anonymous):

thx

jimthompson5910 (jim_thompson5910):

yw